OpenStudy (anonymous):

Sovle x y y=4x-2 y=3x+5 lightly explain. Thanx

OpenStudy (anonymous):

subtract 2 equns. to eliminate y 0 = x - 7 x = 7 plug x = 7 in first equn, y =4(7) - 2 y = 26